WAC Notes August 16 2020
From PREX Wiki
|HOW TOs for shift crew|
|All Expert Contacts|
- 1 August 16 2020
- 2 Non-Production Runs and Bad Runs
- 3 Issues
- 4 Updates
August 16 2020
Daily Meeting BlueJeans
Production Run List - Slug 190 - IHWP OUT - Flip-RIGHT
Continued from yeterday's run here
Back to the OLD Moller quads tune - Septum hadn't been cycled after hall work
- Good - Run 7759 - Looks good.
- 100 uA for most of the run.
- First run back with the originally used Moller quad tune set up.
- Noticed the SBS and HES septum magnet power supplies both weren't cycled correctly around noon and cycled them now.
Production Run List - Slug 191 - IHWP OUT - Flip-RIGHT
RHRS Dipole tripped, its own slug
- Good - Run 7760 - Looks good
- Cut out :70k events for large beam motion at the beginning of the run.
- R dipole tripped -> Set RHRS main detectors to local cuts and no-blinding and no-normalization
- Good - Run 7762 - Looks good. 12X BPM jumps, need to cut out >1mm, < -1mm. LHRS only run, R Dipole tripped
- Good - Run 7763 - looks good. LHRS only run
- Good - Run 7764 - Looks good. LHRS only run. Cut out BPM 12X > 0.5
- Good - Run 7765 - beam motion a lot, add cuts to event number, remove 170k - 280k
- Good - Run 7766 - has a jump in 12X midway through... probably OK
Non-Production Runs and Bad Runs
- Junk - Run 7761 - Carbon 1% for compton chicane position optimization
- Junk - Run 7767 - Carbon 1% for compton
- Running with FFB off
- Moller quads failed and we got a different tune over night. The data seems OK anyway
- Run 7761 for Compton tuning on C-1%, needed vertical table move
- BMOD tune seems OK.
- I found another bug in code - was always using 4aX BPM even in 1X case for corrections... now fixed
- It appears 1X 15746 BMOD analysis is the best for the last few days of running. Needs re-analysis to pick up pedestal fix and get the full run lists all analyzed again.